OpenStudy (anonymous):

MATHS HELP NEEDED EXAM TOMORROW! 50 MULITPLE CHOICE PRACTICE QUESTIONS :)

OpenStudy (anonymous):

if a seller incerease the price of his product by p%, by what percentage does he need to lower the new price to get back to the original price? a) \[p \div (1-p \div100)\] b) p c)\[100p \div(100-p)\] d) \[100p \div(100+p)\]

OpenStudy (anonymous):

step by step working please :)

OpenStudy (shubhamsrg):

let original price be x after inc, new price = x + (p% of x) = x+ px/100 = x( 1+ p/100) now you want to get back to original price i.e. x % dec = (x-x(1+ p/100)) / (x(1 + p/100) *100 =(1 - 1 + p/100) / ( 1 + p/100) *100 = (p/100+p) * 100 hope it helps../

OpenStudy (anonymous):

thanks shubhamsrg, my percentage was wrong, i just wrote it as p% :)

OpenStudy (shubhamsrg):

nevermind,,glad to help ^_^

OpenStudy (anonymous):

in order to paint a cube, 1 liter paint is needed. the cube is cut into 27 identical cubes. how much paint is needed to paint the 27 little cubes? a)1L b)3L c)9L d)27L

OpenStudy (anonymous):

@Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

1 big cube of length a has area 6a² to be painted 27 small cubes of length .... have total area of 27 x .... to be painted. Then direct proportionality should help you find the answer is 3 L.

OpenStudy (anonymous):

@Vincent-Lyon.Fr thank you!